# The expected kinematic matter dipole is robust against source evolution

### v2

@@ Abstract
Recent measurements using catalogues of quasars and radio galaxies have shown that the dipole anisotropy in the large-scale distribution of matter is about twice as large as is expected in the standard $\Lambda$CDM model, indeed in any cosmology based on the Friedman-Lemaître-Robertson-Walker (FLRW) metric. This expectation is based on the kinematic interpretation of the dipole anisotropy of the cosmic microwave background, i.e. as arising due to our local peculiar velocity. The effect of aberration and Doppler boosting on the projected number counts on the sky of cosmologically distant objects in a flux-limited catalogue can then be calculated and confronted with observations. This fundamental consistency test of FLRW models proposed by Ellis & Baldwin (1984) was revisited by Dalang & Bonvin (2022) who argued that redshift evolution of the sources can significantly affect the expected matter dipole. In this note we demonstrate that the Ellis & Baldwin test is in fact robust to such effects, hence the $>5\sigma$ dipole anomaly uncovered by Secrest et al. (2021, 2022) remains an outstanding challenge to the $\Lambda$CDM model.
@@
Large-scale Structure – Cosmological Principle
@@ 1 Introduction
We are not in fact comoving observers; we infer this from the observation of a dipole anisotropy in the Cosmic Microwave Background (CMB) which is interpreted as due to our peculiar velocity $\beta\sim 10^{-3}$ with respect to the rest frame in which the CMB is isotropic (Stewart & Sciama, 1967; Peebles & Wilkinson, 1968). This is called the ‘CMB frame’, and the CP implies that the large-scale distribution of matter should also be isotropic in this frame. Hence in order to interpret cosmological data in the standard $\Lambda$CDM framework, observables measured in the heliocentric frame are boosted to the CMB frame. This has shown good agreement with a wide range of data, moreover direct tests of homogeneity have so far supported the CP. For example counts-in-cells of quasars (e.g. Laurent et al., 2016) indicate homogeneity is reached on scales larger than a few hundred Mpc, although worryingly our local peculiar ‘bulk flow’ shows no signs of dying out even on such scales (e.g. Watkins et al., 2023). With the exploration of cosmic structure on increasingly large scales, tests of the CP underlying $\Lambda$CDM have become increasingly important.
@@ 2.1 The magnification bias $x$
Figure 1: Sketch of the integral number counts (black curve) indicating that the index $\tilde{x}$ is to be measured at the limiting flux $S_{*}$ (dashed line) by fitting a power-law (blue line) in its neighbourhood. Only sources with flux density $S$ in a narrow (shaded) region of fractional width $\mathcal{O}(\beta)$ around $S_{*}$ contribute to the dipole amplitude, so the slope at higher or lower flux density is irrelevant.
We follow the notation in DB22, but here, for completeness, state the general and implicit dependences $x(r)=x(r,S_{*})$ and $f(r)=f(r,S_{*})$. In each shell of $r$ we find a distribution of sources with observed spectral flux density $S>S_{*}$,
@@ 2.2 The spectral index $\alpha$
While DB22 propose an effective value $\alpha_{\rm eff}$ as defined in Eq. (3) we reiterate that the relevant quantity should not be averaged over all sources, but only over those that lie close enough to the threshold $S_{*}$ such that they actually contribute to the dipole amplitude.
Rather than integral source counts we require knowledge of differential source counts that can be evaluated at a given flux density $S$. The integral and differential source counts are related by
@@
This can be used to define a second normalised distribution function over comoving distance, $\tilde{f}(r)=\tilde{f}(r,S_{*})$,
@@
where we have used Eqs. (8) and again (4). Now $\tilde{\alpha}$ is computed as the average spectral index of sources at the threshold $S_{*}$,
@@
The difference between $\tilde{\alpha}$ and $\alpha_{\rm eff}$ is illustrated by the marginalised distributions shown in Fig. 2: the 2-D distribution ${\rm d}N/{\rm d}S{\rm d}\alpha(S,\alpha)$ returns different 1-D distributions of $\alpha$, if averaged over all $S>S_{*}$, or only in the neighbourhood of $S_{*}$. This highlights the importance of correlations between $S$ and $\alpha$ in a given catalogue; these were in fact accounted for in the mock sky simulations of Secrest et al. (2021). While the area around the flux limit within which sources contribute to the dipole amplitude depends on $\alpha$ (as indicated by the white shaded region in the figure), in practice such corrections are negligible. Lastly, if $x(r)$ is a constant, then $\tilde{\alpha}=\alpha_{\rm eff}$ trivially.
Figure 2: Left panel: Illustrative sketch of the two-dimensional distribution of $S$ and $\alpha$. The spectral index $\tilde{\alpha}$ is to be measured locally at $S_{*}$ (vertical dashed line). As in Fig. 1, only sources with flux density $S$ in a region (grey shaded) close to $S_{*}$ contribute to the dipole amplitude, so the $\alpha$ for other sources are irrelevant. Right panel: Marginalised distributions of $\alpha$ when averaged over all values of $S>S_{*}$ (blue curve) or just over those in the vicinity of $S_{*}$ (black curve) leading to potentially different average values (horizontal dashed lines).
@@ Acknowledgements
It is a pleasure to thank Charles Dalang, Harry Desmond, Nathan Secrest, and especially Subir Sarkar, for useful discussions on this matter, and for comments on an initial draft. I am also grateful for many helpful remarks by and exchanges with Camille Bonvin, Chris Clarkson, Ruth Durrer, Pedro Ferreira, and Roy Maartens following the first circulation of this manuscript.

### v3

@@ Abstract
Recent measurements using catalogues of quasars and radio galaxies have shown that the dipole anisotropy in the large-scale distribution of matter is about twice as large as is expected in the standard $\Lambda$CDM model, indeed in any cosmology based on the Friedman-Lemaître-Robertson-Walker (FLRW) metric. This expectation is based on the kinematic interpretation of the dipole anisotropy of the cosmic microwave background, i.e. as arising due to our local peculiar velocity. The effect of aberration and Doppler boosting on the projected number counts on the sky of cosmologically distant objects in a flux-limited catalogue can then be calculated and confronted with observations. This fundamental consistency test of FLRW models proposed by Ellis&Baldwin in 1984 was revisited recently arguing that redshift evolution of the sources can significantly affect the expected matter dipole. In this note we demonstrate that the Ellis&Baldwin test is in fact robust to such effects, hence the $>5\sigma$ dipole anomaly uncovered recently remains an outstanding challenge to the $\Lambda$CDM model.
@@
large-scale structure of Universe – cosmology: observations – cosmology: theory
@@ 1 Introduction
We are not in fact comoving observers; we infer this from the observation of a dipole anisotropy in the Cosmic Microwave Background (CMB) which is interpreted as due to our peculiar velocity $\beta\,=v/c\sim 10^{-3}$ with respect to the rest frame in which the CMB is isotropic (Stewart & Sciama, 1967; Peebles & Wilkinson, 1968). This is called the ‘CMB frame’, and the CP implies that the large-scale distribution of matter should also be isotropic in this frame. Hence in order to interpret cosmological data in the standard $\Lambda$CDM framework, observables measured in the heliocentric frame are boosted to the CMB frame. This has shown good agreement with a wide range of data, moreover direct tests of homogeneity have so far supported the CP. For example counts-in-cells of quasars (e.g. Laurent et al., 2016) indicate homogeneity is reached on scales larger than a few hundred Mpc, although worryingly our local peculiar ‘bulk flow’ shows no signs of dying out even on such scales (e.g. Watkins et al., 2023). With the exploration of cosmic structure on increasingly large scales, tests of the CP underlying $\Lambda$CDM have become increasingly important.
@@ 2.1 The magnification bias $x$
Figure 1: Sketch of the integral number counts (black curve) indicating that the index $\tilde{x}$ is to be measured at the limiting flux $S_{*}$ (dashed line) by fitting a power-law (blue line) in its neighbourhood. Only sources with flux density $S$ in a narrow (shaded) region of fractional width $\mathcal{O}(\beta)$ around $S_{*}$ (here chosen to be $S_{*}^{\rm true}$) contribute to the dipole amplitude, so the slope at higher or lower flux density is irrelevant. In practice, measurement of the slope $\tilde{x}$ must be performed at $S_{*}^{\rm true}$ in bins small enough to resolve the shaded region.
We follow the notation in DB22, but here, for completeness, state the general and implicit dependences $x(r)=x(r,S_{*})$ and $f(r)=f(r,S>S_{*})$. In each shell of $r$ we find a distribution of sources with observed spectral flux density $S>S_{*}$,
@@ 2.2 The spectral index $\alpha$
While DB22 propose an effective value $\alpha_{\rm eff}$ as defined in Eq. (3) we reiterate that the relevant quantity should not be averaged over all sources, but only over those that lie close enough to the threshold $S_{*}$ such that they actually contribute to the dipole amplitude. In other words, since a source $i$ adds to the kinematic dipole only if its flux density $S_{i}$ lies within $S_{*}\pm(1+\alpha_{i})\beta S_{*}$ (depending on its location on the sky), we must not also average over those sources’ spectral indices $\alpha_{i}$ that are not included in this interval. We therefore require a corresponding formalism that restricts the average value $\alpha_{\rm eff}$ of Eq. (3) which runs over all $S>S_{*}$ as per the definition of $f(r)$, to on that only includes sources with $S\sim S_{*}$.
To this effect, rather than integral source counts we require knowledge of differential source counts that can be evaluated at a given flux density $S$. The integral and differential source counts are related by
@@
This can be used to define a second normalised distribution function over comoving distance, $\tilde{f}(r)=\tilde{f}(r,S_{*})$, only of sources with flux densities $S=S_{*}$,
@@
where we have used Eqs. (8) and again (4). Thus $\tilde{\alpha}$ is computed as the average spectral index of sources at the threshold $S_{*}$ as,
@@
The difference between $\tilde{\alpha}$ and $\alpha_{\rm eff}$ is illustrated by the marginalised distributions shown in Fig. 2: the 2-D distribution ${\rm d}N/{\rm d}S{\rm d}\alpha(S,\alpha)$ returns different 1-D distributions of $\alpha$, if averaged over all $S>S_{*}$, or only in the neighbourhood of $S_{*}$. This highlights the importance of correlations between $S$ and $\alpha$ in a given catalogue; these were in fact accounted for in the mock sky simulations of Secrest et al. (2021). While the area around the flux limit within which sources contribute to the dipole amplitude depends on $\alpha$ (as indicated by the white shaded region in the figure, and noted at the beginning of this subsection), in practice such corrections are negligible. Lastly, if $x(r)$ is a constant, then $\tilde{\alpha}=\alpha_{\rm eff}$ trivially.
Figure 2: Left panel: Illustrative sketch of the two-dimensional distribution of $S$ and $\alpha$. The spectral index $\tilde{\alpha}$ is to be measured locally at $S_{*}$ (vertical dashed line). As in Fig. 1, only sources with flux density $S$ in a region (grey shaded) close to $S_{*}$ contribute to the dipole amplitude, so the $\alpha$ for other sources are irrelevant (The white shaded region indicates the true functional dependence on $\alpha$ of the relevant region, see main text for details). The white dotted lines are for illustrative purpose only and denote iso-potential lines of the distribution. Right panel: Marginalised distributions of $\alpha$ when averaged over all values of $S>S_{*}$ (blue curve) or just over those in the vicinity of $S_{*}$ (black curve) leading to potentially different average values (horizontal dashed lines).
@@ Acknowledgements
It is a pleasure to thank Charles Dalang, Harry Desmond, Nathan Secrest, and especially Subir Sarkar, for useful discussions on this matter, and for comments on an initial draft. I am also grateful for many helpful remarks by and exchanges with Camille Bonvin, Chris Clarkson, Ruth Durrer, Pedro Ferreira, and Roy Maartens following the first circulation of this manuscript. Finally, I thank the referee for their constructive remarks and suggestions to improve the paper.
@@
For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
Data Availability
No new data were generated or analysed in support of this research.